a ball is dropped from a building and simultaneously another ball is projected upward with some velocity.describe the change in relative velocities of the ball as a function of time (IIT JEE) drops of water fall at regular intervals from the roof of a building of height 16m. the first drop strikes the ground at the same moment as the fifth drop leaves the roof . find the distance between the 2 nd and 3 rd drops. (AIEEE-2007)

  1. a ball is dropped from a building and simultaneously another ball is projected upward with some velocity.describe  the change in relative velocities of the ball as a function of time (IIT JEE)
  2. drops of water fall at regular intervals from the roof of a building of height 16m. the first drop strikes the ground at the same moment as the fifth drop leaves the roof . find the distance between the 2nd and 3rd drops. (AIEEE-2007)

Grade:11

2 Answers

Samyak Sharma
5 years ago
If you know the basics then you can solve it 
 
Remember that there are 4 drops in air simultaneously  
So take ratio like
 
1x+3x+5x+7x= 16
I.e. X=1
 
Distance between 2nd and 3rd drop is 5-3 = 2m
The RockStar
3 years ago
1A.... Initial Velocity of freely falling ball = 0
 
Initial velocity of projected ball = u 
 Relative velocity = u+0 =u
Final velocity of freely falling ball
 
👉 V=u+at, V1=0+gt, V1 = gt
 
Final Velocity (V2) of projected ball
 
👉V=u+at, v=u-gt   
 Relative velocity = gt + u - gt (because they are in opposite direction relative velocities are added)  = u
 
Relative velocity is independent on Time....
 
Change in Relative velocity = u - u =0 😄